OpenStudy (anonymous):

what is pi to 4 decimal places?

sammixboo (sammixboo):

What do you already know of pi?

OpenStudy (anonymous):

That It's on your pic?

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

And we need to know 4 places to the left or to the right? @shayiq21926

sammixboo (sammixboo):

Reply then write @sammixboo so I know you are finished =3 \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\)

jagr2713 (jagr2713):

\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\]

OpenStudy (michele_laino):

I remember that: \[\pi = 3.14159...\]

OpenStudy (anonymous):

only michele-laino was a bit of help